# Entanglement phases and phase transitions in monitored free fermion systems of localization

### v2

@@ Abstract
In recent years, the presence of local potentials has significantly enriched and diversified the entanglement patterns in monitored free fermion systems. In our approach, we employ the stochastic Schrödinger equation to simulate a one-dimensional spinless fermion system under continuous measurement and local potentials. By averaging the steady-state entanglement entropy over many quantum trajectories, we investigate its dependence on measurement and localization parameters. We used a phenomenological model to interpret the numerical results, show that the introduction of local potentials does not destroy the universality class of the entanglement phase transition, and the phase boundary is jointly characterized by the measurement process and the localization mechanism. This work offers a new perspective on the characterization of the entanglement phase boundary arising from the combined effects of measurement and localization, and provides criteria for detecting this novel phase transition in cold atom systems, trapped ions, and quantum dot arrays.
@@ I Introduction
Classical paradigms of nonequilibrium quantum physics-such as thermalization under the eigenstate thermalization hypothesis [1, 2, 3, 4, 5] and its counterpoint in the quantum Zeno effect [6, 7, 8, 9, 10], offer a striking tension: while coherent dynamics generically scramble information and build volume-law entanglement, frequent observations can freeze evolution by continually projecting local degrees of freedom. This tension naturally leads to the notion of a measurement-driven reorganization of quantum states: as the rate or strength of monitoring is increased, the balance between entanglement production and removal can qualitatively change, suggesting a sharp transition in the long-time entanglement structure. This phenomenon is now understood as a measurement-induced phase transition (MIPT) [11, 12, 13, 14, 15, 16, 17], in which entanglement scaling morphs from volume law to area law due to the backaction of measurements.
A broad body of work has established both the ubiquity and diversity of MIPTs across platforms and protocols [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. In random quantum circuits [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], interleaving unitary gates with stochastic measurements yields a finite-threshold transition. Related transitions appear under Hamiltonian dynamics with projective or continuous monitoring [51, 52, 53, 54, 55], and persist for both projective and weak measurements. Across these settings, the steady-state phases (volume, logarithmic/sub-volume, and area) and their critical properties depend on symmetries [48, 58, 59, 19], dimensionality  [60], and the microscopic structure of measurements [37, 61, 62, 63, 64, 65], revealing a rich taxonomy of measurement-driven universality classes.
@@
Against this backdrop, we ask two connected questions at the intersection of measurement and localization: (I) Will a change in the localization mechanism alter the critical universality class of the MIPT phase diagram, or do the two cases follow the same critical laws? (II) is it possible to formulate a unified understanding of entanglement phase transitions arising from the interplay between distinct localization mechanisms and measurements? Addressing these questions disentangles universal from non-universal ingredients in the measurement-localization competition and yields operational criteria for experimental tests across platforms ranging from cold atomic systems [52, 91, 92, 51, 53, 93] and trapped ions [94, 95] to quantum dot arrays [96].
@@
Figure 2: Phase diagrams and phase boundaries extracted from the effective central charge. (a) Phase diagram under joint tuning of the measurement strength and SP strength; (b) Phase diagram under joint tuning of the measurement strength and the QPP strength. Colors indicate the effective central charge $\bar{c}_{\mathrm{eff}}$, extracted from the finite-size scaling of half-chain entanglement entropy: green corresponds to large $\bar{c}_{\mathrm{eff}}$ values (log-law-like regime), while brown corresponds to $\bar{c}_{\mathrm{eff}}\rightarrow 0$ (area-law regime). Dashed curves denote the semi-analytical expressions for phase boundaries derived by Eq. (12). Black solid dots mark four numerically fitted phase transition points and the grey region corresponds to the case of a very weak localization potential, where we tend to revert to the results obtained from the Keldysh field theory [60]; therefore, it is not shown in the diagram. In the subsequent finite-size scaling analysis, four representative points on each phase diagram (marked in red, green, blue, and purple, respectively) are selected to characterize the entanglement scaling behavior at the phase boundary of that point.
We consider a spinless one-dimensional fermions model with a localization, subject to continuous measurements:
@@
$P_{j}$ represents the strength of potential at lattice site $j$:
$P_{j}=\begin{cases}\Delta j/L,&\text{SP},\\ V\cos\left(\dfrac{2\pi j}{\tau}+\theta\right),&\text{QPP},\end{cases}$ (2)
where $J$ is the strength of the nearest-neighbor hopping term, $c_{j}$ and $c^{\dagger}_{j}$ are the annihilation and creation operators of the spinless fermion at site $j$. Here, $\Delta$ denotes the SP strength, $L$ is the number of lattice sites in the system, and the parameters $V$, AA, and $B$ represent, respectively, the QPP strength, the golden mean $(\sqrt{5}+1)/2$, and the phase shift of the potential. We initialize the system in a Néel state, that is, $\psi(t=0)=$ $\prod_{i=1}^{L/2}c_{2i-1}^{\dagger}|\mathrm{vac}\rangle$, where $|\mathrm{vac}\rangle$ is the vacuum state, and consider its evolution under open boundary conditions (OBCs). We perform a quantum measurement uniformly for all sites with a finite measurement amplitude (free quency) $\gamma$:
@@
$+\sum_{j=1}^{L}\left[\frac{c_{j}^{\dagger}c_{j}|\psi(t)\rangle}{\sqrt{\langle\psi(t)|n_{j}|\psi(t)\rangle}}-|\psi(t)\rangle\right]dW_{j}(t),$ (3)
where $dW_{j}(t)$ takes 0 or 1 obeying the site-independent Poisson process, that is, $\left\langle\left\langle dW_{j}(t)\right\rangle\right\rangle=\gamma dt$. Here, $\langle\langle\ldots\rangle\rangle$ is the average noise. As a result of the measurement process, there are many trajectories of wave functions starting from the fixed initial state. Hence, we need to take an average over many trajectories to examine the statistical properties of any observable. The main physical quantity that we address is entanglement entropy. Divide the system into two subsystems $A$ and $B$, which have sizes $\ell$ and $L-\ell(\ell\leq L/2)$, respectively ($L$ represents the system size). Then, the entanglement entropy is defined as:
@@ III.1 Phase boundaries
where $\rho=|\psi(t)\rangle\langle\psi(t)|,h_{AB}:=H_{AB}/\left\|H_{AB}\right\|(\|\ldots\|$ is the operator norm), and $\mathbf{I}_{B}$ is the identity operator for subsystem $B$. The Hamiltonian $H_{AB}$ denotes the boundary interaction between subsystems $A$ and $B$ :
@@
where $h_{i,j}$ is an interaction operator acting on sites $i$ and $j$. Of interest is the case where $\ell=L/2$. In our study of a fermionic system without long-range interactions, the entanglement entropy growth rate is finite and takes a constant value, denoted as $G_{0}$. Since measurements and local potentials both act to disentangle the system, even when small, the original volume-law entanglement entropy growth $S\sim G_{0}t$ is transformed, under their combined effect, into a logarithmic form $S\sim\log(t)$, and, upon further increase, crosses over to an area-law scaling $S\sim t^{0}$. In our study, the phase boundary corresponds to the critical line separating the log-law region from the area-law region. Therefore, we find that at the phase boundary the entanglement entropy growth follows $S\sim C_{0}\log(t)$, representing the lower bound of the log-law region and the upper bound of the area-law region (see Appendix. A), where $C_{0}$ is a small constant. Assuming that under the influence of measurements and local potentials the initial entanglement growth rate is renormalized to $\Gamma(\gamma,P)G_{0},$ then at criticality one can write:
@@
$\Gamma(\gamma)$ represents the entanglement suppression rate induced by the measurement strength $\gamma$. It quantifies the degree to which entanglement growth is suppressed per unit time under continuous measurement interventions. When $\gamma$ is sufficiently large, entanglement growth can be strongly inhibited or even fully halted. To analyze this process quantitatively, we decompose the system’s evolution into individual quantum trajectories. Each trajectory corresponds to a specific measurement history, and the dynamics are governed by an effective non-Hermitian Hamiltonian:
@@
where $H$ is the original Hermitian Hamiltonian, $\gamma$ is the measurement rate, and $\mathcal{L}_{u}$ is the Lindblad operator. The non-Hermitian term captures the non-unitary dissipative effects introduced by measurements. Under this non-Hermitian evolution, Under the trajectory-weighted average, the initial entanglement-entropy growth rate of the system in the presence of measurements is renormalized to (please refer to Appendix.A for details):
@@
As shown in Eq. (6), under purely unitary evolution, entanglement growth is driven by the boundary coupling between subsystems $A$ and $B$. The entanglement growth rate is fundamentally determined by the Hamiltonian term $H_{AB}$ that connects $A$ and $B$. In the presence of a local potential, the spatial localization of particle states significantly suppresses the boundary coupling between regions, leading to a pronounced reduction in the entanglement growth rate [101]. Corresponding to the distinct localization properties of the wave functions in the SP and the QPP systems, the entanglement growth rate is modified by the local potentials as:
@@
Eq.(12) provides a phenomenological envelope-function expression [73], and in the Anderson case [102], the measurement further introduces a super-exponential envelope that strongly suppresses the coefficient $C_{0}$. To validate Eq. (12), we first map the effective central charge $\bar{c}_{\mathrm{eff}}$ over the parameter space spanned by the measurement strength $\gamma$ and the local potential strength (either Stark, $\Delta$, or quasi-periodic, $V$), thereby characterizing the scaling regimes of the entanglement entropy. As shown in Fig. 2, the phase diagrams under varying measurement and potential strengths are presented for the SP and QPP systems. The colormap encodes $\bar{c}_{\mathrm{eff}}$ extracted from half-chain entanglement entropy; the horizontal axis is the measurement rate $\gamma$, and the vertical axis is the potential strength $\Delta$ or $V$. The diagrams clearly separate area-law and log-law regimes: regions with large $\bar{c}_{\mathrm{eff}}$ correspond to log-law growth (an entanglement-spreading phase), whereas regions with $\bar{c}_{\mathrm{eff}}\approx 0$ exhibit suppressed entanglement.
@@ III.2 Phases and phase transitions
We next follow the finite-size flow of the effective central charge $\bar{c}_{\mathrm{eff}}$ under OBCs. For each size $L$ we fit the entanglement profile to Eq. (13) and obtain a single $\bar{c}_{\mathrm{eff}}$ subjected to different measurement intensities and local potentials, as illustrated in Fig. 4. Near criticality, the half-chain entanglement entropy exhibits logarithmic scaling consistent with non-unitary conformal field theory [104, 105]:
@@
where $\ell$ is the subsystem size, $\bar{c}_{\mathrm{eff}}$ is the effective central charge, and $s_{0}$ is a non-universal constant. In the thermodynamic limit, $\bar{c}_{\mathrm{eff}}\to 0$ in the area-law phase, while it remains finite in the logarithmic phase, thus serving as a diagnostic for the transition [93, 106]. As shown in Fig. 4(a,b), when the local potential is fixed at a small value ($\Delta=0.1$ or $V=0.3$), increasing the measurement strength $\gamma$ causes $\bar{c}_{\mathrm{eff}}$ curves for different system sizes to gradually intersect, indicating a transition near $\gamma\approx 0.3$. The insets strongly corroborate this behavior: as $\gamma$ increases, the entanglement entropy transitions from rapid growth to saturation, signaling a crossover from log-law to area-law scaling. Similarly, in Fig. 4(c,d), when the measurement strength is fixed at a low value ($\gamma=0.1$), increasing the local potential strength leads to similar crossing behavior among different system sizes, with transition points occurring around $\Delta\approx 0.3$ and $V\approx 1.8$, respectively. The insets in Fig. 4 clearly demonstrate this effect: the entanglement entropy drops from steep growth to a nearly constant value as localization increases, again signifying a transition from log-law to area-law behavior. These results collectively indicate that both measurement strength and local potential can induce a phase transition in the system. In both cases, entanglement growth is suppressed, although the mechanisms and suppression profiles differ between SP and QPP. These observations reveal how both measurement and localization reduce boundary coupling and suppress entanglement growth. The distinct functional behaviors of SP and QPP reinforce the validity of $\bar{c}_{\mathrm{eff}}$ as a transition indicator.
@@
where $\bar{S}_{A}$, $\bar{S}_{B}$, and $\bar{S}_{AB}$ denote the entanglement entropies of regions $A$, $B$, and $A\cup B$, respectively. Mutual information has proven to be a sensitive diagnostic of MIPT in a variety of systems [39, 53], due to its ability to capture nonlocal correlations. In our study, we adopt the same set of parameters used for computing $\bar{c}_{\mathrm{eff}}$. As shown in Fig. 5(a,b), when the local potential is set to a small value ($\Delta=0.1$ or $V=0.3$), increasing the measurement strength $\gamma$ leads to a crossing in mutual information curves across different system sizes, signaling the onset of a transition. Similarly, in Fig. 5(c,d), fixing the measurement strength and gradually increasing the potential strength ($\Delta$ or $V$) also induces such crossings, indicative of a phase boundary. To further understand the origin of these behaviors, we compute the connected correlation function defined as
@@
which reduces to $\bar{C}_{\ell}=|\langle\hat{c}^{\dagger}_{L/2}\hat{c}_{L/2+\ell}\rangle|^{2}$ for the Slater-determinant state $|\psi(t)\rangle$. As shown in the insets of Fig. 5, both increasing measurement strength and local potential lead to qualitatively similar changes in the correlation structure. In the low measurement or weak localization regime, $\bar{C}_{\ell}$ decays algebraically, $\bar{C}_{\ell}\sim\ell^{-\eta}$, indicating long-range correlations and rapid entanglement growth. In the strong measurement or deep localized regime, $\bar{C}_{\ell}$ decays exponentially, $\bar{C}_{\ell}\sim\exp(-\ell/\xi)$, suggesting short-range correlations and saturation of entanglement entropy consistent with area-law behavior. This sharp decay in the correlation function confirms a transition from an entangled phase with extended quantum coherence to a localized phase dominated by measurement or potential-induced suppression. Together with mutual information crossings, these results demonstrate that both measurement and disorder drive a phase transition that limits entanglement growth, with distinct suppression mechanisms depending on the type of local potential applied.
Table 1: Data collapse parameters for the effective central charge results.
$\Delta/V$ $\gamma(\Delta)/\gamma(V)$
Data 0.1/0.3 0.16/0.8 0.23/1.3 0.3/1.8 0.16/0.1 0.23/0.18 0.28/0.24 0.3/0.3
$\gamma_{c}$ or $\Delta_{c}/V_{c}$ 0.3/0.3 0.28/0.24 0.23/0.18 0.16/0.1 0.1/0.3 0.16/0.8 0.23/1.3 0.3/1.8
$\nu$ 3.8/2.7 4.1/2.9 5.7/3.3 6.5/3.8 3.3/2.6 4.1/2.9 4.9/3.2 6.5/3.6
$\alpha$ 5.9/4.1 6.2/4.3 6.5 /4.5 7.3 /4.9 4.9 /3.2 6.5/4.1 7.4/4.8 8.9/5.3

### v4

# Entanglement phases and phase transitions in the monitored free fermion systems of localization

@@ Abstract
The interplay between quantum measurement and localization significantly alters the spreading of quantum information. In this work, we investigate monitored free-fermion chains with localized potentials. With the aid of the quantum trajectory method and finite size analysis, we numerically reveal that the Berezinskii–Kosterlitz–Thouless transition, which emerges in monitored free-fermion chains, is robust in the presence of localization. To understand this, we construct a phase diagram that diverges at a boundary described by entanglement propagation. We find that the monitored system wtih Stark-localized decays quickly under small measurement strength, whereas the Anderson-localized system decays more slowly. By incorporating localization effects, this work advances our understanding of how measurement competes with the coherent spreading of quantum information. Our findings can potentially be realized in cold atom systems, trapped ions, and quantum dot arrays.
@@ I Introduction
Classical paradigms of nonequilibrium quantum physics-such as thermalization under the eigenstate thermalization hypothesis [31, 100, 86, 26, 55] and its counterpoint in the quantum Zeno effect [79, 50, 34, 56, 33], offer a striking tension: while coherent dynamics generically scramble information and build volume-law entanglement, frequent observations can freeze evolution by continually projecting local degrees of freedom. This tension naturally leads to the notion of a measurement-driven reorganization of quantum states: as the rate or strength of monitoring is increased, the balance between entanglement production and removal can qualitatively change, suggesting a sharp transition in the long-time entanglement structure. This phenomenon is now understood as a measurement-induced phase transition (MIPT) [36, 65, 69, 68, 111, 124, 48, 82], in which entanglement scaling morphs from volume law to area law due to the backaction of measurements.
A broad body of work has established both the ubiquity and diversity of MIPTs across platforms and protocols [37, 43, 106, 94, 96, 91, 107, 95, 81, 59, 18, 15, 110, 116, 52, 53, 62, 3, 66, 22, 97, 67, 41, 42, 51, 10, 23, 101, 35, 109, 61, 89, 49, 104, 40, 38, 88, 73, 102, 76, 5, 12, 83, 60, 24, 103, 30, 121]. In random quantum circuits [66, 22, 97, 67, 41, 42, 51, 10, 23, 101, 35, 109, 61, 89, 49], interleaving unitary gates with stochastic measurements yields a finite-threshold transition. Related transitions appear under Hamiltonian dynamics with projective or continuous monitoring [104, 40, 38, 88, 73], and persist for both projective and indirect measurements. Across these settings, the steady-state phases (volume, logarithmic/sub-volume, and area) and their critical properties depend on symmetries [61, 5, 12, 43], dimensionality  [83], and the microscopic structure of measurements [22, 60, 24, 103, 30, 121], revealing a rich taxonomy of measurement-driven universality classes.
@@
Previous studies in disordered [102] and quasi-periodic systems [76] have revealed the characteristics of measurement-induced phase transitions, with their results mainly emphasizing the persistence of the BKT universality class and the regulation of phase laws by Anderson localization potentials. In contrast to these studies based on Anderson localization, we choose the Stark localization system as a stronger localization platform, taking this as an opportunity to further explore the competition between measurement and different localization mechanisms, and attempt to construct a phase diagram that diverges at a boundary characterized by entanglement propagation under distinct localization mechanisms.
Against this backdrop, we ask two connected questions at the intersection of measurement and localization: (I) Will a change in the localization mechanism alter the critical universality class of the MIPT phase diagram, or do the two cases follow the same critical laws? (II) Is it possible to establish a unified phase diagram to characterize entanglement phase transitions arising from the interplay between different localization mechanisms and measurements? Addressing these questions disentangles universal from non-universal ingredients in the measurement-localization competition and yields operational criteria for experimental tests across platforms ranging from cold atomic systems [40, 25, 32, 104, 38, 6] and trapped ions [80, 4] to quantum dot arrays [54].
@@
Figure 2: Phase diagrams and phase boundaries extracted from the effective central charge. (a) Phase diagram under joint tuning of the measurement strength and SP strength; (b) Phase diagram under joint tuning of the measurement strength and the QPP strength. Colors indicate the effective central charge $\bar{c}_{\mathrm{eff}}$, extracted from the finite-size scaling of half-chain entanglement entropy: green corresponds to large $\bar{c}_{\mathrm{eff}}$ values (log-law-like regime), while brown corresponds to $\bar{c}_{\mathrm{eff}}\rightarrow 0$ (area-law regime). Dashed curves denote the semi-analytical expressions for phase boundaries derived by Eq. (12). Black solid dots mark four phase transition points obtained through finite-size scaling analysis and data collapse. The grey region corresponds to the case of a very weak localization potential, where we tend to revert to the results obtained from the Keldysh field theory [83]; therefore, it is not shown in the diagram. In the subsequent finite-size scaling analysis, four representative points on each phase diagram (marked in red, green, blue, and purple, respectively) are selected to characterize the entanglement scaling behavior at the phase boundary of that point.
We consider a spinless one-dimensional fermions model of localization, subject to continuous measurements:
@@
where $P_{j}$ represents the strength of potential at lattice site $j$:
$P_{j}=\begin{cases}\Delta j/L,&\text{SP},\\ V\cos\left(\dfrac{2\pi j}{\tau}+\theta\right),&\text{QPP},\\ \varepsilon_{j},&\text{Anderson disorder},\end{cases}$ (2)
with $J=1$ is the strength of the nearest-neighbor hopping term, $c_{j}$ and $c^{\dagger}_{j}$ are the annihilation and creation operators of the spinless fermion at site $j$. Here, $\Delta$ denotes the SP strength, $L$ is the number of lattice sites in the system, and the parameters $V$, $\tau$, and $\theta$ represent, respectively, the QPP strength, the golden mean $(\sqrt{5}+1)/2$, and the phase shift of the potential. The on-site potential $\varepsilon_{j}$ is an independent random variable drawn from the uniform distribution $\left[-{W},{W}\right]$, where $W$ is the disorder strength. We initialize the system in a Néel state, that is, $\psi(t=0)=$ $\prod_{i=1}^{L/2}c_{2i-1}^{\dagger}|\mathrm{vac}\rangle$, where $|\mathrm{vac}\rangle$ is the vacuum state, and consider its evolution under open boundary conditions (OBCs).We perform continuous measurements uniformly on all lattice sites with a finite measurement rate $\gamma$, where quantum-jump events occur according to the standard quantum-jump trajectory formalism[28, 29, 113, 117] . The corresponding stochastic Schrödinger equation is given by [1, 6, 38, 114]
@@
$\quad+\sum_{j=1}^{L}\left[\frac{c_{j}^{\dagger}c_{j}|\psi(t)\rangle}{\sqrt{\langle\psi(t)|n_{j}|\psi(t)\rangle}}-|\psi(t)\rangle\right]dW_{j}(t),$ (3)
where $n_{j}=c_{j}^{\dagger}c_{j}$ is the local particle number operator being monitored. The stochastic variables $dW_{j}(t)$ take values 0 or $1$ according to independent Poisson processes with mean $\langle\!\langle dW_{j}(t)\rangle\!\rangle=\gamma\langle n_{j}\rangle dt,$ and $\langle\!\langle\cdots\rangle\!\rangle$ denotes the noise average. In this formulation, the Born probabilities are encoded in the normalization factor $\sqrt{\langle\psi(t)|n_{j}|\psi(t)\rangle}$, which ensures that the probability of each measurement outcome is consistent with quantum mechanics, while the Poisson increments $dW_{j}(t)$ determine whether a measurement event occurs at site $j$ during each infinitesimal time step. As a result, the system evolves along many possible quantum trajectories starting from the same initial state, and statistical properties of observables are obtained by averaging over a large ensemble of such trajectories without any postselect. We use 500 independent trajectories in our numerical simulations. The main physical quantity that we address is entanglement entropy. Divide the system into two subsystems $A$ and $B$, which have sizes $\ell$ and $L-\ell(\ell\leq L/2)$, respectively ($L$ represents the system size). Then, the entanglement entropy is defined as:
@@ III.1 Phase boundaries
where $\rho=|\psi(t)\rangle\langle\psi(t)|,h_{AB}:=H_{AB}/\left\|H_{AB}\right\|(\|\ldots\|$ is the operator norm), and $\mathbf{I}_{B}$ is the identity operator for subsystem $B$. The Hamiltonian $H_{AB}$ denotes the boundary interaction between subsystems $A$ and $B$ [108] :
@@
where $h_{i,j}$ is an interaction operator acting on sites $i$ and $j$. Of interest is the case where $\ell=L/2$. In our study of a fermionic system without long-range interactions, the entanglement entropy growth rate is finite and takes a constant value, denoted as $G_{0}$. Since measurements and localized potentials both act to disentangle the system, even when small, the original volume-law entanglement entropy growth $S\sim G_{0}t$ is transformed, under their combined effect, into a logarithmic form $S\sim\log(t)$, and, upon further increase, crosses over to an area-law scaling $S\sim t^{0}$. In our study, the phase boundary corresponds to the critical line separating the log-law region from the area-law region. Therefore, we find that at the phase boundary the entanglement entropy growth follows $S\sim C_{0}\log(t)$, representing the lower bound of the log-law region and the upper bound of the area-law region (see Appendix. A), where $C_{0}$ is a small constant. Assuming that under the influence of measurements and localized potentials the initial entanglement growth rate is renormalized to $\Gamma(\gamma,P)G_{0},$ then at criticality one can write:
@@
$\Gamma(\gamma)$ represents the entanglement suppression rate induced by the measurement strength $\gamma$. It quantifies the degree to which entanglement growth is suppressed per unit time under continuous measurement interventions. When $\gamma$ is sufficiently large, entanglement growth can be strongly inhibited or even fully halted. To analyze this process quantitatively, we decompose the system’s evolution into individual quantum trajectories. Each trajectory corresponds to a specific measurement history, and the dynamics are governed by an effective non-Hermitian Hamiltonian [6]:
@@
where $H$ is the original Hermitian Hamiltonian, $\gamma$ is the measurement rate, and $\mathcal{L}_{u}$ is the Lindblad operator. The non-Hermitian [119, 118, 71, 84, 57, 92, 123, 122, 98, 99, 93, 85, 115, 120] term captures the non-unitary dissipative effects introduced by measurements. Under this non-Hermitian evolution, Under the trajectory-weighted average, the initial entanglement-entropy growth rate of the system in the presence of measurements is renormalized to (please refer to Appendix.A for details):
@@
As shown in Eq. (6), under purely unitary evolution, entanglement growth is driven by the boundary coupling between subsystems $A$ and $B$. The entanglement growth rate is fundamentally determined by the Hamiltonian term $H_{AB}$ that connects $A$ and $B$. In the presence of a localized potential, the spatial localization of particle states significantly suppresses the boundary coupling between regions, leading to a pronounced reduction in the entanglement growth rate [11]. Corresponding to the distinct localization properties of the wave functions in the SP and the QPP systems, the entanglement growth rate is modified by the localized potentials as:
@@
Eq.(12) provides a phenomenological envelope-function expression [39], and in the Anderson case [58], the measurement further introduces a super-exponential envelope that strongly suppresses the coefficient $C_{0}$. To validate Eq. (12), we first map the effective central charge $\bar{c}_{\mathrm{eff}}$ over the parameter space spanned by the measurement strength $\gamma$ and the localized potential strength (either Stark, $\Delta$, or quasi-periodic, $V$), thereby characterizing the scaling regimes of the entanglement entropy. As shown in Fig. 2, the phase diagrams under varying measurement and potential strengths are presented for the SP and QPP systems. The colormap encodes $\bar{c}_{\mathrm{eff}}$ extracted from half-chain entanglement entropy; the horizontal axis is the measurement rate $\gamma$, and the vertical axis is the potential strength $\Delta$ or $V$. The diagrams clearly separate area-law and log-law regimes: regions with large $\bar{c}_{\mathrm{eff}}$ correspond to log-law growth (an entanglement-spreading phase), whereas regions with $\bar{c}_{\mathrm{eff}}\approx 0$ exhibit suppressed entanglement.
@@ III.2 Phases and phase transitions
We next follow the finite-size flow of the effective central charge $\bar{c}_{\mathrm{eff}}$ under OBCs. For each size $L$ we fit the entanglement profile to Eq. (13) and obtain a single $\bar{c}_{\mathrm{eff}}$ subjected to different measurement intensities and localized potentials, as illustrated in Fig. 4. Near criticality, the half-chain entanglement entropy exhibits logarithmic scaling consistent with non-unitary conformal field theory [20, 21]:
@@
where $\ell$ is the subsystem size, $\bar{c}_{\mathrm{eff}}$ is the effective central charge, and $s_{0}$ is a non-universal constant. In the thermodynamic limit, $\bar{c}_{\mathrm{eff}}\to 0$ in the area-law phase, while it remains finite in the logarithmic phase, thus serving as a diagnostic for the transition [6, 17]. As shown in Fig. 4(a,b), when the localized potential is fixed at a small value ($\Delta=0.1$ or $V=0.3$), increasing the measurement strength $\gamma$ causes $\bar{c}_{\mathrm{eff}}$ curves for different system sizes to gradually intersect, indicating a transition near $\gamma\approx 0.3$. The insets strongly corroborate this behavior: as $\gamma$ increases, the entanglement entropy transitions from rapid growth to saturation, signaling a crossover from log-law to area-law scaling. Similarly, in Fig. 4(c,d), when the measurement strength is fixed at a low value ($\gamma=0.1$), increasing the localized potential strength leads to similar crossing behavior among different system sizes, with transition points occurring around $\Delta\approx 0.3$ and $V\approx 1.8$, respectively. The insets in Fig. 4 clearly demonstrate this effect: the entanglement entropy drops from steep growth to a nearly constant value as localization increases, again signifying a transition from log-law to area-law behavior. These results collectively indicate that both measurement strength and localized potential can induce a phase transition in the system. In both cases, entanglement growth is suppressed, although the mechanisms and suppression profiles differ between SP and QPP. These observations reveal how both measurement and localization reduce boundary coupling and suppress entanglement growth. The distinct functional behaviors of SP and QPP reinforce the validity of $\bar{c}_{\mathrm{eff}}$ as a transition indicator.
@@
where $\bar{S}_{A}$, $\bar{S}_{B}$, and $\bar{S}_{AB}$ denote the entanglement entropies of regions $A$, $B$, and $A\cup B$, respectively. Mutual information has proven to be a sensitive diagnostic of MIPT in a variety of systems [67, 38], due to its ability to capture nonlocal correlations. In our study, we adopt the same set of parameters used for computing $\bar{c}_{\mathrm{eff}}$. As shown in Fig. 5(a,b), when the localized potential is set to a small value ($\Delta=0.1$ or $V=0.3$), increasing the measurement strength $\gamma$ leads to a crossing in mutual information curves across different system sizes, signaling the onset of a transition. Similarly, in Fig. 5(c,d), fixing the measurement strength and gradually increasing the potential strength ($\Delta$ or $V$) also induces such crossings, indicative of a phase boundary. To further understand the origin of these behaviors, we compute the connected correlation function defined as
@@
which reduces to $\bar{C}_{\ell}=|\langle\hat{c}^{\dagger}_{L/2}\hat{c}_{L/2+\ell}\rangle|^{2}$ for the Slater-determinant state $|\psi(t)\rangle$. As shown in the insets of Fig. 5, both increasing measurement strength and localized potential lead to qualitatively similar changes in the correlation structure. In the low measurement or weak localization regime, $\bar{C}_{\ell}$ decays algebraically, $\bar{C}_{\ell}\sim\ell^{-\eta}$, indicating long-range correlations and rapid entanglement growth. In the strong measurement or deep localized regime, $\bar{C}_{\ell}$ decays exponentially, $\bar{C}_{\ell}\sim\exp(-\ell/\xi)$, suggesting short-range correlations and saturation of entanglement entropy consistent with area-law behavior. This sharp decay in the correlation function confirms a transition from an entangled phase with extended quantum coherence to a localized phase dominated by measurement or potential-induced suppression. Together with mutual information crossings, these results demonstrate that both measurement and disorder drive a phase transition that limits entanglement growth, with distinct suppression mechanisms depending on the type of localized potential applied.